OpenStudy (anonymous):

What is the approximate hydronium ion concentration and hydroxide ion concentration in bleach? Which is higher?

OpenStudy (anonymous):

Will give medals

OpenStudy (anonymous):

Bleach is basic so it has a higher hydroxide ion concentration.

OpenStudy (anonymous):

The pH of bleach is roughly 12; recall that the pH and pOH must sum to 14, so we find pOH = 14 - 12 = 2. To determine concentration, recall:$$pOH=-\log[OH^-]\\ [OH]^-=10^{-pOH}=10^{-2}=1/100$$